OpenStudy (amonoconnor):

I have, yes... Another question regarding the evaluation of the integral of the following: (sin(x^(1/2))/(x^(1/2)))dx, from x=(pi^2)/4 to pi^2 I am setting out to evaluate this integral as I have the rest of those I've done today- through u-substitution. So... 1. Can/should I set u = x^(1/2) ? 2. If not, then what is the appropriate/best thing to use as "u"? Any and all help is greatly appreciated! :)

OpenStudy (kmeis002):

Yes, that u will work. Its best to use u-sub if you see a function and a form of its derivative. For yours \[ u = \sqrt{x} \\\frac{du}{dx} = \frac{1}{2\sqrt{x}}\] which we see in the integral, except for the constant

OpenStudy (amonoconnor):

Alright, that makes sense... I'm going to work through the problem:) will you be online/ monitoring this question for a while??? :D

OpenStudy (kmeis002):

I will for another half hour or so

OpenStudy (amonoconnor):

Okay, so I have this so far, but am kind of confused as to the next step: u = x^(1/2) du = 1/2(x^(-1/2))dx And... integ.(sin(u))(u^-1) Also... dx = du/(1/2(x^-1/2)) = (du/1)((2(x^(-1/2)))/1) = (2x^(-1/2))du So is the following also correct, considering "dx"? 2x^(-1/2)du = (2u)du (?)

OpenStudy (amonoconnor):

Making the integral: (sin(u))(u^-1)(2u)du ... Does that "break rules"?

OpenStudy (kmeis002):

Hmm, here's what I would get step by step \[u = \sqrt{x} \\ \frac{du}{dx} = \frac{1}{2\sqrt{x}} \to du=\frac{1}{2\sqrt{x}} dx \to 2du=\frac{1}{\sqrt{x}} dx \] Now we can replace our \[ \frac{1}{\sqrt{x}} dx\] to obtain \[\int \frac{\sin{\sqrt{x}}}{sqrt{x}} dx \to \int \sin{u}\: 2du \to2\int \sin{u}\: du \]

OpenStudy (amonoconnor):

Ahhh... Okay, I follow. :) why didn't I see that?! Lol...

OpenStudy (amonoconnor):

Thank you! I see how that works out now :)

OpenStudy (kmeis002):

Not a problem, glad I could help

OpenStudy (amonoconnor):

Oh, sorry, I know I closed the question, but after the step you posted above, when I remove the integral sign, and am taking the anti derivative, do I use the chain rule on theta (u), making it: 2(-cosu)u^2 ?

OpenStudy (amonoconnor):

@kmeis002

OpenStudy (amonoconnor):

:)

OpenStudy (kmeis002):

No need, now we are integrating with respect to u (du). The u-substitution is reversing the chain rule, so it takes care of it. The anti-differentive is just -cos(u) + C, then just plug in the sqrt(x)

OpenStudy (amonoconnor):

Thank you!! :)

OpenStudy (kmeis002):

not a problem.